OpenStudy (anonymous):

2(4x-4)+2(3x+2)=360 how do i do this step by step?

OpenStudy (anonymous):

8x - 8 + 6x + 6 = 360 ok?

OpenStudy (anonymous):

Solve for x: 2 (3 x+2)+2 (4 x-4) = 360

OpenStudy (anonymous):

it would be 6x+6+8x+8

OpenStudy (anonymous):

Factor the polynomial, 4 x-4.

OpenStudy (anonymous):

im just afraid I'm doing it wrong with combining the like terms

OpenStudy (anonymous):

No,No!you are doing it wrong!

OpenStudy (anonymous):

Solve for x: 2 (3 x+2)+2 (4 x-4) = 360 Factor the polynomial, 4 x-4. Factor 4 from the polynomial 4 x-4: 2 (3 x+2)+2×4 (x-1) = 360 Multiply 2 and 4 together. 2×4 = 8: 2 (3 x+2)+8 (x-1) = 360 Distribute 2 over 3 x+2. 2 (3 x+2) = 6 x+4: 6 x+4+8 (x-1) = 360 Distribute 8 over x-1. 8 (x-1) = 8 x-8: 8 x-8+6 x+4 = 360 Group like terms in 8 x+6 x-8+4. Grouping like terms, 8 x+6 x-8+4 = (6 x+8 x)+(4-8): (6 x+8 x)+(4-8) = 360 Add like terms in 6 x+8 x. 6 x+8 x = 14 x: 14 x+(4-8) = 360 Evaluate 4-8. 4-8 = -4: 14 x+-4 = 360 Factor the polynomial, 14 x-4. Factor 2 from the polynomial 14 x-4: 2 (7 x-2) = 360 Divide both sides by a constant to simplify the equation. Divide both sides of 2 (7 x-2) = 360 by 2: (2 (7 x-2))/2 = 360/2 Any nonzero number divided by itself is one. 2/2 = 1: 7 x-2 = 360/2 Reduce 360/2 to lowest terms. The gcd of 360 and 2 is 2, so 360/2 = (2×180)/(2×1) = 2/2×180 = 180: 7 x-2 = 180 Isolate terms with x to the left hand side. Add 2 to both sides: 7 x+(2-2) = 180+2 Look for two terms that sum to zero. 2-2 = 0: 7 x = 180+2 Evaluate 180+2. 180+2 = 182: 7 x = 182 Solve for x. Divide both sides of 7 x = 182 by 7: (7 x)/7 = 182/7 Any nonzero number divided by itself is one. 7/7 = 1: x = 182/7 Reduce 182/7 to lowest terms. The gcd of 182 and 7 is 7, so 182/7 = (7×26)/(7×1) = 7/7×26 = 26: Answer: | | x = 26

OpenStudy (anonymous):

6x+4 my bad

OpenStudy (anonymous):

whoa lol

OpenStudy (anonymous):

@zkrup , Don't Give him the answer!

OpenStudy (anonymous):

Show him the steps!

OpenStudy (anonymous):

I don't want the answer I just need help with the step

OpenStudy (anonymous):

OK,I will help you...

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

2(4x-4)+2(3x+2) = 8x - 8 + 6x + 2

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

do we subtract 6x on each side?

OpenStudy (anonymous):

@PFEH.1999 Do you not see all those steps?

OpenStudy (anonymous):

and it is: 8(x - 1) + 6(x + 2)

OpenStudy (anonymous):

@zkrup !but you gave him the answer!

OpenStudy (anonymous):

hes just helpin

OpenStudy (anonymous):

@PFEH.1999 But I showed him how to get it.

OpenStudy (anonymous):

ok,ok I know!he helped well!

OpenStudy (anonymous):

but he shouldn't give the answer...

OpenStudy (anonymous):

omg can we stop this lol

OpenStudy (anonymous):

ik im talking about you

OpenStudy (anonymous):

did you understand this? 8(x - 1) + 6(x + 2)

OpenStudy (anonymous):

so do I add the like terms now?

OpenStudy (anonymous):

no :/

OpenStudy (anonymous):

ok,so i should show you another way...

OpenStudy (anonymous):

is it 8 times x and 1 + 6 times x +2

OpenStudy (anonymous):

plese forgot that :) 8x - 8 + 6x + 2 = 360 --> 8x + 6x = 8 -2 + 360

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

8 - 2 + 360 = ? say...

OpenStudy (anonymous):

366

OpenStudy (anonymous):

ok,and this one? 8x + 6x = ?

OpenStudy (anonymous):

14x

OpenStudy (anonymous):

oh sorry,It was 8 - 6 + 360 not 8 -2 + 360 ... and x = 362 / 14

OpenStudy (anonymous):

really I though there were re steps?

OpenStudy (anonymous):

i'll do it again just to make sure I got it and thanks a lot!

OpenStudy (anonymous):

ok do it...perhaps i made some mistakes :)

OpenStudy (anonymous):

I think we forgot some last steps but its ok i'll finsh it gtg though thanks you very much

OpenStudy (anonymous):

2(4x-4)+2(3x+2)=360 Distribute the number outside the parenthesis to the number inside the parenthesis. Basically, you multiply the number outside the parenthesis by EACH number inside the parenthesis. 8x-8+6x+4=360 Now, Combine like terms. 8x+6x & -8+4 14x-4=360 +4 +4 Separate the variable. So you add +4 on both sides to get variable alone. 14x=364 Isolate the variable. Divide by 14 on both sides, so that in the end... It's JUST x. x=26

OpenStudy (anonymous):

@FilthyMcNasty , solved it correct :)